OpenStudy (anonymous):

how would you find d^2y/dx^2 of: y=sin^2(2x)?

OpenStudy (math&ing001):

\[\frac{ d ^{2}y }{ dx ^{2} }=\frac{ d }{ dx }(\frac{ dy }{ dx })\] Do you know how to derivate?

OpenStudy (anonymous):

yep, i first tried to find the 1st derivative of sin^2 2x and got 4sin2xcos2x but can't seem to find the derivative of that :/

OpenStudy (math&ing001):

You can find the derivative of 4sin2xcos2x by using this rule : d/dx(f*g)=f*d/dx(g)+g*d/dx(f)

OpenStudy (cwrw238):

yes math is right this is called the product rule used to differentiate the product of 2 functions so you need to differentiate 4sin2x and cos2x - can you do that?

OpenStudy (anonymous):

i thought so but i keep getting the wrong answer :( would you do 4sin2x x -2sin2x + cos2x x 8cos2x?

OpenStudy (cwrw238):

thats correct

OpenStudy (cwrw238):

it can be simplified

OpenStudy (anonymous):

thanks, but do you know how you would get this to 8cos4x? which is the answer I'm meant to end up with :)

OpenStudy (cwrw238):

ok - i think you use the identity cos2x = cos^2x - sin^2x if we replace x by 2x we get cos4x = cos^2 2x - sin^2 2x

OpenStudy (anonymous):

CALCULATOR BRAH 2ND DEGREE SOMETHING LIKE THAT AND PUT SIN AND TYPE THAT IN BOOOOM

OpenStudy (cwrw238):

and we can simplify your result to 8( cos^2 2x - sin^2 2x) do you see the connection?

OpenStudy (cwrw238):

8( cos^2 2x - sin^2 2x) = 8 cos 4x

OpenStudy (anonymous):

yes, thank you very much! :D